OpenStudy (kkutie7):

A driver needs to stop for a red light and slows down at a rate of 4m/s^2. it takes 100m to stop. (a) Find the time required to stop. (b) Find the initial velocity. (c) Plot x(t), v(t), and a(t). I need to be walk through this one.

OpenStudy (aaronq):

Are you familiar with the kinematic equations of motion?

OpenStudy (aaronq):

\(\sf v_f=v_i+a\Delta t\) \(\sf x_f=x_i+v_o\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\) \(\sf v_f^2=v_i^2+2a_x\Delta x\)

OpenStudy (kkutie7):

I'm also not allowed to use any other equation except these: \[v^{\rightarrow}=\frac{dr^{\rightarrow}}{dt}\] \[a^{\rightarrow}=\frac{dv^{\rightarrow}}{dt}\] \[v^{\rightarrow}=v _{0}^{\rightarrow}+at^{\rightarrow}\] \[r^{\rightarrow}=r _{0}^{\rightarrow}+v _{0}^{\rightarrow}t+\frac{1}{2}a^{\rightarrow}t^{2}\]

OpenStudy (kkutie7):

ok so you used the \[r^{\rightarrow}\] one

OpenStudy (aaronq):

okay, no problem. Use the last equation you posted to find the time.

OpenStudy (aaronq):

wait, you dont know the initial velocity.

OpenStudy (kkutie7):

nope I have to figure that out too

OpenStudy (aaronq):

so you can't use \(\sf v_f^2=v_i^2+2a_x\Delta x\)?

OpenStudy (kkutie7):

no =(

OpenStudy (kkutie7):

can we use the fact that the ending velocity is zero?

OpenStudy (aaronq):

yeah, we need to use that fact, but without using that last equation i posted, all the equations have 2 unknown variables.

OpenStudy (aaronq):

I guess we can use \(\sf v_f=v_i+a\Delta t\) solve for \(\sf v_i\), and use that into \(\sf x_f=x_i+v_o\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\)

OpenStudy (kkutie7):

yeah I see that. so I have a and I have the r?

OpenStudy (aaronq):

then our only unknown is the time.

OpenStudy (kkutie7):

ok so for the first one will look like this \[0=v^{\rightarrow}t+4t\]

OpenStudy (aaronq):

theres no t with the initial velocity and depending which way you set up the axis, the acceleration will take a negative or positive sign. \(\sf 0=v_i-4\Delta t\)

OpenStudy (kkutie7):

my mistake

OpenStudy (aaronq):

no worries, now just sub that into the other equation, and solve for t

OpenStudy (kkutie7):

ok let me do that real quick

OpenStudy (aaronq):

okay, it should look like this \(\sf 100=4\Delta t^2 -\dfrac{1}{2}4(\Delta t)^2\)

OpenStudy (kkutie7):

what happened to the xi

OpenStudy (aaronq):

well normally, you wanna set up your coordinate system so that the initial value is zero, so \(x_i=0\) and it's just omitted.

OpenStudy (kkutie7):

shouldn't it be more like this? \[(4\Delta t^{2})\Delta t\]

OpenStudy (kkutie7):

ok that makes sense

OpenStudy (aaronq):

no it's fine the way it is

OpenStudy (aaronq):

because \(v_i=4t\) substituing that in gives what i wrote earlier

OpenStudy (kkutie7):

no I get that but in the equation I have its vi*delta t

OpenStudy (aaronq):

sorry about leaving suddenly, something came up and i had to go. Hope you got the question.